# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v2

@@ 5 End-to-end security: Biological invariant
If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong. These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin [14], or if the pump insertion site hits a vein and delivers insulin directly into the blood stream [62]. Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user. For the purposes of the following example, suppose the end-to-end check is that the measured glucose levels are always within 30mg/dl of the predicted level; we refer to this as the implementation invariant that is enforced in an end-to-end way across the system:
$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<30\mathsf{mg/dl}\quad(1)$
Naively, one might look at the $30\mathsf{mg/dl}$ on the right-hand side and think that the user’s true glucose levels are within $30\mathsf{mg}$ of the actual glucose levels. But this would be wrong! The problem is that the measured glucose level is subject to measurement error; the state-of-the-art Dexcom G7 CGM has a mean absolute relative difference of 8.2% from the true glucose level in the body [29]. For the sake of example, suppose that the measurement error on the CGM reading is around 5%; we refer to this as the CGM invariant:
$|g_{\mathsf{measured}}-g_{\mathsf{actual}}|<5\mathsf{mg/dl}\quad(2)$
and from this we can derive the true conclusion on the user’s glucose levels, which is what we refer to as the biological invariant. We stress that this invariant involves the parameter $g_{\mathsf{actual}}$ that is not available in software; it instead ensures a property (which we are able to formally prove) about the real level of glucose in the biological system. That statement we prove in the software is actually the following, not (1):
$|g_{\mathsf{measured}}-g_{\mathsf{predicted}}|<35\mathsf{mg/dl}\quad(3).$
When we run BioKernel on a real human, we don’t want to assume safety with respect to the measured glucose levels (invariant (1)); we actually want to ensure safety with respect to the real glucose levels in the human that is using the system (invariant (3)). Failing to distinguish these cases could be a severe software bug, and would jeopardize the end-to-end safety that we aim to provide by failing to account for CGM error. In the rest of the section, we describe the invariants that we enforce in the implementation (including invariant (1) as a simple case) and what additional invariants we prove in our formal verification development (including invariant (3)) below.
Implementation invariant.
@@
In the implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs. If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be. We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detectability of significant glycemic deviations above this level [37, 34].
@@
Verification.
In the verification development (described in Section 7), we prove that the biological invariant holds, given error bounds set by the CGM manufacturer.
First, in the ported Dafny code, we locate the portions which enforce implementation invariants described above, including invariant (1). The code includes the two parameters $g_{\mathsf{measured}}$ and $g_{\mathsf{actual}}$, as these are available to the software implementation. Second, we model actual glucose by putting error bounds on our CGM readings, in the CGM invariant (2). This CGM invariant is inserted as a precondition on all functions that operate on CGM values. Third, we track the individual’s actual glucose level using a ghost variable in Dafny. Ghost variables are variables that are only available to the prover, these do not exist in the real software system. Fourth and lastly, we express as a postcondition the biological invariant (4), which asserts that the absolute difference between the theoretical model glucose and the ground-truth actual glucose cannot exceed a threshold of 35 mg/dl per hour. When the code is compiled, the Dafny verifier proves that the biological invariant holds on all inputs to the verified function.
Evaluation results.
Our evaluation shows that the biological invariant is effective at keeping humans safe when they face drastic changes in insulin sensitivity (Section 8.3), and that our thresholds of 30 mg/dl per hour and two hours were effective for the human that used GlucOS to manage their T1D without any false positives (Section 9.3).
Novelty.
We beleive that having a parameter real biological state of the system is an important modeling decision; and it diverges, for example, from previous work which only models the biological data in simulation [71, 72]. Traditional software verification techniques like model checking [36] and deductive verification [27] have been explored in the past for hybrid system verification [26] that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process. Also, researchers have used real-world invariants for security in the context of autonomous vehicles [18], but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
6 Last line of defense: Humans
In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system. While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance.
When an individual’s glucose levels are out-of-range, they need to intervene to bring them back in range. Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered. To recover from hypoglycemia, people need to eat sugar. Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
Figure 6: GlucOS predicts hypoglycemia and notifies the individual before it happens so they can proactively prevent it.
We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 6), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack. Alerts for hypoglycemia are not new. All CGM software comes with non-predictive hypoglycemia alerts by default [2] and glucose prediction algorithms have been well covered in the literature [65, 70, 11]. Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
@@ 6 Last line of defense: Humans
Our finding has the potential to have broader implications for security in healthcare systems. Often, security systems favor uniformity and consistency [23, 28], but when managing people’s health we found that allowing people to balance alert fatigue and security increased engagement and potentially increased protection, based on each individual’s needs.
7 Formal verification
@@
To provide additional assurance that our implementaiton is correct, we port our Swift code to Dafny and use Hoare logic. As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example [19]. Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent. We port updated code back to Swift manually because Dafny does not include code generation for Swift.
@@
These efforts paid off: we found and fixed nine bugs in our original implementation (Appendix D); in addition to proving various invariants in the code for our insulin delivery path and security mechanisms. However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms. For a full list of the invariants that we define and prove, see Appendix E.
8 Evaluation
8.1 How effective is GlucOS’s algorithmic security mechanism?
We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator [7]. We use simulation since this evaluation involves injecting dangerous amounts of insulin.
@@ 8.1 How effective is GlucOS’s algorithmic security mechanism?
To be effective, GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D. Accordingly, we first run simulations with a predictive ML model inspired from prior research [52], without incorporating our security mechanism to set a baseline of an algorithm that is efficient at managing T1D. Then, in the same simulation scenarios, we introduce an order of magnitude difference (increase and decrease) to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms. We then repeat both sets of simulations while incorporating our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
@@
The goal of T1D management is to maintain glucose levels that are in a safe, target range (70-180 mg/dl) for the majority of the time – an outcome that has been linked to positive long-term health outcomes [53]. Glucose levels 54-69 mg/dl (level 1 hypoglycemia) or less than 54 mg/dl (level 2 hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels 181-250 mg/dl (level 1 hyperglycemia) or greater than 250 mg/dl (level 2 hyperglycemia) increase risks for long-term vascular damage. The American Diabetes Association (ADA) recommends that individuals spend at least 70% of their time in range [6, 24] in order to remain healthy.
8.1.1 Without algorithmic security mechanism
@@
Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training. Figure 7 shows the average proportion of time spent in range (TIR) by virtual humans from different age groups (adults, adolescents and children) when employing the predictive ML model. We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model. These results provide a baseline for an algorithm that is effective in managing T1D.
@@
Figure 7: Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
@@
Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms. To simulate attacks, we intentionally introduce an order of magnitude difference to the amount of insulin predicted by the baseline predictive ML model. In one experiment, we dose one-tenth of the amount of insulin and in another experiment, we dose ten times the amount of insulin computed by the predictive ML model. We run these malicious ML algorithms within the simulator on each virtual human on the same scenarios as the previous experiment.
@@
Figure 8 and Figure 9 show the CGM readings on three randomly chosen virtual humans on a particular scenario for the experiments where we dose less and dose more respectively. When dosing less, we see that all of them spend most of their time above Level 1 hyperglycemia. When dosing more, all individuals fully go below Level 2 hypoglycemia, which would have killed them.
Figure 8: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin. The CGM readings staying mostly above the Level 1 hyperglycemic range.
@@
Figure 9: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin. The CGM readings staying below the Level 2 hypoglycemic range would have likely been fatal to all three individuals.
Across all virtual humans, we see an average drop in TIR from 88.44% with correct dosing to 23.32% when dosing less and an average drop in TIR from 88.44% to 0% when dosing more. Correspondingly, when dosing less, we observe an average increase in the time in hyperglycemia from 3.77% with correct dosing to 65.05%. Similarly, when dosing more, we observe an average increase in the time in hypoglycemia from 0% to 100%.
These results further motivate the need for security mechanisms when providing individuals with the flexibility to use any algorithm to manage T1D.
8.1.2 With algorithmic security mechanism
@@
Impact on incorrect dosing. We repeat the experiments from the previous section with the same scenarios. We run the same models for incorrect dosing, but also run our reactive safe model (Section 3) to set bounds on the amount of insulin the malicious algorithms can inject.
@@
Figure 10 and Figure 11 show CGM readings on the same three virtual humans on the same scenarios from the experiment in the previous section. With both types of incorrect dosing, we see that all of them spend over 70% of their time in range. These results show that GlucOS protects individuals even in presence of malicious incorrect dosing. On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 73.89% when employing safety while dosing less and a drop in TIR from 88.44% to 78.22% when dosing more. While the times spent in range is less than that of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing an active attack.
Figure 10: Variation in CGM readings for three virtual individuals when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin. On average, the CGM readings stay in range over 70% of the time.
@@
Figure 11: Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin. On average, CGM readings stay in range over 70% of the time.
@@
Impact on effective benign algorithms. We evaluate the impact of GlucOS’s algorithmic security mechanism on benign algorithms that are effective in managing T1D such as the baseline predictive ML model. We repeat the experiment of running the predictive ML model on all virtual humans in the same random scenarios while maintaining the same parameters for the clamps and reactive model used to evaluate incorrect dosing.
@@
Figure 12: We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands. Our driver security mechanism detected all manipulations.
8.2 Does GlucOS ensure pump operation while overcoming malicious pump drivers?
We first evaluate GlucOS’s driver security mechanism in terms of its impact on benign pump drivers that accurately relay pump communications. We then evaluate the impact on detecting additional, modified, or dropped communication from a malicious pump driver. Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver. We do not use the simulator for this evaluation since it does not incorporate a pump driver.
Impact on malicious pump drivers. As shown in Figure 12, we attached an Omnipod DASH insulin pump to a stuffed animal. We then modified the source code of the OmniBLE pump driver to intentionally add, drop and modify insulin commands. We detected all manipulated pump commands showing that our driver security mechanism has 0 false negatives in detecting manipulated commands.
Impact on benign pump drivers. For one week, we logged messages from the BioKernel to unmodified OmniBLE as well as messages sent from OmniBLE to the pump on an individual running GlucOS (See Section 9). We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanisms do not incur any false positives.

### v3

@@ 5. End-to-end security: Biological invariant
When implementing security systems for humans, even the perfect algorithmic and driver security mechanisms are unable to handle all situations due to changes in human physiology. A key novelty in our overall approach is our introduction of an end-to-end check that ensures the BioKernel’s glucose calculations against observations of the user’s metabolic state. If there is a large mismatch between our calculations and what we observe, then it means that our calculations are wrong. These deviations can happen from drastic changes in the user’s glucose absorption, like during exercise when the body uses glucose without needing insulin (borghouts2000exercise, ), or if the pump insertion site hits a vein and delivers insulin directly into the blood stream (steyn2023accelerated, ). Regardless of the cause, the end-to-end check detects these conditions and prevents harm to the user. We call this safety check that accounts for changing human physiology a biological invariant.
@@
In our implementation, we check that changes in glucose levels that we observe from CGM readings and the individual’s insulin delivery history are consistent with our theoretical calculations of how much the person’s glucose should change given these same inputs. If their glucose is dropping by more than our theoretical calculations show, then they are more sensitive to insulin then our system believes they should be. We set a threshold of 30 mg/dl per hour, for the maximum allowable divergence based on clinical evaluations showing detection of significant glycemic deviations above this level (30mg_1, ; 30mg_2, ).
@@
6. Last line of defense: Humans
In this section, we describe our design and implementation for incorporating humans as a critical component of our multi-layered security system. While our algorithmic, system-level, and biological invariant defenses form the primary barriers against attacks, humans serve as the last line of defense, capable of mitigating attacks that manifest as glycemic imbalance. We discuss our most interesting findings around preferences and the architectural iterations we make to our design based on our real-world experiences.
6.1. Predictive alerts for humans
When an individual’s glucose levels are not within the range of healthy values, they need to intervene to bring them back in range. Low glucose, or hypoglycemia, occurs from too much insulin, and there is no way to remove insulin that the automated insulin delivery system has already delivered. To recover from hypoglycemia, people need to eat sugar. Given that people with T1D are already active participants in managing glycemic imbalance, we focus on how we can use this fact to our advantage to create a more comprehensive security system.
Figure 5. GlucOS predicts hypoglycemia and notifies the user before it happens so they can proactively prevent it.
We use predictive yet deterministic mobile notifications to alert people when they are likely to experience hypoglycemia before they reach it (Figure 5), as hypoglycemia presents the most immediate health risk and could be an indicator of an attack. Alerts for hypoglycemia are not new. All CGM software comes with non-predictive hypoglycemia alerts by default (dexcom_alerts, ) and glucose prediction algorithms have been well covered in the literature (Josep2020, ; Zhang2021, ; arora2023multivariate, ). Our contribution is using predictive alerts before someone reaches hypoglycemia as a security defense.
@@ 6.1. Predictive alerts for humans
Our finding has the potential to have broader implications for security in healthcare systems. Often, security systems favor uniformity and consistency (DHILLON2016656, ; FURNELL20165, ), but when managing people’s health we found that allowing people to balance alert fatigue and security increased engagement and potentially increased protection, based on each individual’s needs.
6.2. Architectural implications
One interesting design iteration we made was moving the alerting logic into the BioKernel itself. In our first design, consistent with our separation principles, the alerting logic resided in a separate app. This alerting app learned the state of the system using our cloud-based event logging system. However, one of our participants went on a canoeing trip and was disconnected from the internet, so they could not get alerts. Ideally iOS would provide inter-process communication mechanisms to pass data from the BioKernel directly to the alerting app, but iOS does not have the appropriate APIs for this style of communication. Thus, we moved the alerting logic into the BioKernel, violating our principle of separation, but enabling our users to get alerts even when they are disconnected from the internet.
Moving the alerting logic into the BioKernel did have security implications. It adds complexity to our trusted computing base, but we carefully manage this tradeoff. We minimize the added complexity by using traditional, well-understood algorithms for prediction rather than more complex ML models and reuse existing functionality within the BioKernel when possible to keep the implementation simple. This approach allowed us to maintain a high level of confidence in the security and correctness of our alerting system while still providing the necessary functionality.
7. Formal verification
@@
To provide additional assurance that our implementation is correct, we port our Swift code to Dafny and use Hoare logic. As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example (chong2020code, ). Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent. We port updated code back to Swift manually because Dafny does not include code generation for Swift.
@@
These efforts paid off: we found and fixed nine bugs in our original implementation; in addition to proving various invariants in the code for our insulin delivery path and security mechanisms. However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms.
Our implementation operates on CGM readings, but in our application of formal methods, our goal is to reason about the individuals actual glucose level, taking into account sensor errors. This raises an interesting challenge for formal verification as it requires modeling the true biological state of the user – a parameter that is both unknown at runtime and distinct from measured parameters that are part of the software implementation. We prove that the biological invariant holds, given error bounds set by the CGM manufacturer (garg2022accuracy, ).
To reason about actual glucose levels formally, we use a ghost variable in Dafny. Ghost variables are variables that are only available to the prover, these do not exist in the real software system. By defining the actual glucose levels using a ghost variable, we can include it in preconditions and postconditions to prove properties about our software relative to actual glucose levels.
We believe that having a parameter for the real biological state of the system is an important modeling decision; and it diverges, for example, from previous work that only models the biological data in simulation (zhou2021data, ; zhou2023knowsafe, ). Traditional software verification techniques like model checking (cps_1, ) and deductive verification (cps_3, ) have been explored in the past for hybrid system verification (cps_2, ) that interact with the human body, but they do not completely account for physiological variability between individuals or model inaccuracies during the insulin delivery process. Also, researchers have used real-world invariants for security in the context of autonomous vehicles (choi2018detecting, ), but these were based on quad-copter physics not human physiology and did not include formal verification of their implementation.
8. Implementation
In our implementation, the BioKernel is an iOS app that consists of 5.1k lines of code in our trusted computing base. It communicates with a Dexcom G6 or G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy, and runs the core closed-loop algorithm. For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state. Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system (Figure 1(c)) that enables separate apps to extend functionality while remaining isolated from the BioKernel.
9. Evaluation
In this section, we evaluate GlucOS’s algorithmic and driver security mechanisms. In the next section (Section 10), we evaluate all components of GlucOS in the real world.
9.1. How effective is GlucOS’s algorithmic security mechanism?
We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator (simglucose, ). GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D. We consider extreme cases of attack in our evaluation. While subtle attacks are also possible, we focus on extreme attacks since they have the most adverse consequences.
@@ 9.1. How effective is GlucOS’s algorithmic security mechanism?
Concreteny, we first run simulations with a predictive ML model inspired from prior research (mosquera2023enabling, ), without incorporating our security mechanism to set a baseline of an effctive algorithm in managing T1D. Then, in the same simulation scenarios, we introduce an order of magnitude increase to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms. We then repeat these experiments with our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
@@
The American Diabetes Association (ADA) recommends that individuals with T1D spend at least 70% of their time in the range (TIR) of 70-180 mg/dl to remain healthy (nathan2014diabetes, ; endocrine_tir, ; journal_glycemic_targets, ). Glucose levels less than 70 mg/dl (hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels greater than 180 mg/dl (hyperglycemia) increase risks for long-term vascular damage.
9.1.1. Without algorithmic security mechanism
@@
Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training. Figure 6 shows the average proportion of time spent in the safe target range (TIR) by virtual humans from different age groups when employing the predictive ML model. We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model. These results provide a baseline for an algorithm that is effective in managing T1D.
@@
Figure 6. Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
@@
Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms. In this section, we discuss an experiment with a malicious algorithm that intentionally doses too much insulin. Concretely, our malicious algorithm doses ten times the amount of insulin computed by the previously described predictive ML model. We ran this algorithm within the simulator on the same virtual human on the same scenarios as the previous experiment.
@@
Figure 7 shows the CGM readings on three randomly chosen virtual humans on a particular scenario when we dose more. The CGM readings of all 3 individuals fully go below the hypogycemic bound, which would have killed them in the real-world.
@@
Figure 7. Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin. The CGM readings staying below the hypoglycemic bound would have likely been fatal to all three individuals.
Across all virtual humans, we see an average drop in TIR from 88.44% to 0% when dosing too much insulin. Correspondingly, we observe an average increase in the time spent in hypoglycemia from 0% to 100%. We also observed similar results with experiments where we intentionally underdosed insulin. In these experiments, all virtual humans spend most of their time above the hyperglycemic bound which wouldn’t have been fatal but would have led to severe health consequences in the real world. These results motivate the need for our algorithmic security mechanism.
9.1.2. With algorithmic security mechanism
@@
Impact on incorrect dosing. We repeat experiments from the previous section with the same scenarios. We run the same malicous model to inject too much insulin, but also run our reactive safe model (Section 3) to bound insulin injection.
@@
Figure 8 show CGM readings on the same three virtual humans on the same scenarios as the experiment in the previous section. We see that all of them spend over 70% of their time in range, showing that GlucOS protects individuals even in presence of malicious overdosing. On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 78.22% when employing safety while overdosing. We also report similar results on experiments with a malicious algorithm that underdoses insulin.
While the times spent in range are less than those of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing active attacks.
@@
Figure 8. Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin. On average, CGM readings stay in range over 70% of the time.
@@
Impact on effective benign algorithms. We evaluate the impact of GlucOS’s algorithmic security mechanism on benign algorithms that are effective in managing T1D such as the baseline predictive ML model. We repeat the experiment of running the predictive ML model on all virtual humans in the same scenarios with the same parameters for the clamps and reactive model used to evaluate incorrect dosing.
@@
Figure 9. We attached an Omnipod DASH insulin pump to a stuffed animal and modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands. Our driver security mechanism detected all manipulations.
9.2. Does GlucOS ensure pump operation while overcoming malicious pump drivers?
Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver.
Impact on malicious pump drivers. We attached an Omnipod DASH insulin pump to a stuffed animal (Figure 9). We then modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands. We detected all manipulated commands thereby showing that our driver security mechanism has 0 false negatives in detecting them.
Impact on benign pump drivers. For one week, we logged messages passed between the BioKernel, the unmodified OmniBLE pump driver and the pump on an individual running GlucOS (Section 10). We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanism does not incur any false positives.